OpenStudy (anonymous):

I really need help with a couple questions will someone please help me please!!!!!!

OpenStudy (shamil98):

post them, please.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

giveme a sec

OpenStudy (anonymous):

what expression is equivalent to this one? –5(6 + 11)

OpenStudy (anonymous):

(-30)+(-55) = -85

OpenStudy (anonymous):

willuhelp me out on like 4 more please

OpenStudy (anonymous):

out of which possibilities? and yeah of course

OpenStudy (anonymous):

Which expression is equivalent to this one? 9(x – 23) A. 9x – 23 B. 9x – 9 · 23 C. 9x · (–23) D. 9 + x – 23

OpenStudy (anonymous):

I used the FOIL system there just so you know how to do it next time

OpenStudy (anonymous):

what is that???

OpenStudy (anonymous):

First Inside Outside Last. Multiplying the outside number by the inside in order in this case so... \[9(x-23) = 9 \times x - 9 \times 23\]

OpenStudy (anonymous):

First Outside Inside Last = FOIL sorry a little misspell the first time

OpenStudy (anonymous):

Have you arrived at an answer yet?